# Gravitational quasinormal modes of black holes in quadratic gravity

### v1

@@ IV.2 Hairy BH background
From (29) we have three nontrivial components. Below we show their schematic form (coefficients depending on the background and $\omega$ are implied):
@@
We then solve the constraint equation together with the $(\theta\varphi)$ component of (29) for $h_{0},\,F_{0}$ and we substitute those into the $(r\varphi)$ component of (29) and the $(r\varphi)$ and $(\theta\varphi)$ components of (30). We end up with the following coupled system of equations for the three perturbation functions $h_{1},\,F_{1},\,F_{2}$
@@
where $\boldsymbol{\Psi}=(\Psi^{(1)},\Psi^{(2)},\Psi^{(3)})\equiv(F_{1},F_{2},h_{1})$, and
@@
where $c$ and $b_{1}$ are the coefficients appearing in the near-horizon expansions of the background solutions (15), (16). By substituting in (61) we can solve for $h_{1}^{(n)},\,F_{1,2}^{(n)}$ with $n\geq 1$ in terms of 3 free parameters, namely $(F_{1}^{(0)},F_{2}^{(0)},h_{1}^{(0)})$.
Figure 5: Real (upper panel) and imaginary (lower panel) parts of the fundamental axial, $\ell=2$ vector modes, as functions of $2M\mu$ in the case of hairy BH background. The solid line correspond to the frequencies of the massless-led and massive-led modes, computed using the DI approach with $n_{\text{max}}=8$. For comparison, we also show the the frequencies of the corresponding modes in the case of the Schwarzschild background (dashed lines). The vertical dashed line denotes the threshold $M=M_{c}$.
@@
For the forward integration the infinity expansions have to be modified in order to incorporate the contribution from the massive and massless mode as well. For QNMs, we look for solutions with $\omega>\mu$ and impose outgoing boundary conditions at infinity, leading to:
@@
where $k=\sqrt{\omega^{2}-\mu^{2}}$ and $x=M(\mu^{2}-2\omega^{2})/(ik)$. After substitution in (61) and solving order by order we determine all coefficients in terms of 3 parameters, namely $(H_{1b}^{(0)},F_{1a}^{(0)},F_{2a}^{(0)})$. To find the QNMs we perform three forward integrations by fixing the horizon parameters $(h_{1}^{(0)},F_{1}^{(0)},F_{1}^{(0)})$, and then demand that the coefficients associated with the ingoing waves at infinity vanish, according to the method explained in Appendix C.
Since the massless and massive perturbations are coupled, we do not have purely massive and purely massless QNM solutions. We find, instead, massless-led solutions, in which the contribution of the massive field is due to the coupling in Eq. (6), and similarly massless-led solutions. Therefore, we have four series of modes: the massless-led vector and tensor modes (which in this case are different from those of GR), and the massive-led vector and tensor modes.
@@
For the vector modes showed in Fig. 5 we notice that within the parameter space we explore, both the decay rate and oscillation frequency increase as the deviation from GR increases, i.e. for smaller values of $\mu$ (larger values of the coupling constant $\alpha$) at a fixed BH mass $M$. In the case of the tensor modes depicted in Fig. 6 the behavior of the massive-led mode is non-monotonic and harder to describe. For clarity, in Figs. 7 and 8, we present the same QNM modes as functions of $p=\mu r_{h}$, and normalizing the frequencies mass with the BH horizon radius rather than with its mass.
At variance with with the case of the Schwarzschild background discussed in the previous subsection, the hairy background is numerical or semi-analytical, and therefore a CF approach for the calculation of QNMs is not as straightforwardly applicable. Thus, we only employ the DI approach. Due to the limitations of the DI method, we do expect numerical errors to be present in this computation. However, the analysis of the Schwarzschild background case gives strong indication that the DI approach with $n_{\text{max}}=8$ is accurate enough to give the QNM frequencies within $\lesssim 1\%$ error for most of the mass range considered.
@@ V Phenomenology
In the region where both Schwarzschild and hairy BHs are radially stable, our analysis did not find any numerical evidence for the instability of either solutions in the axial sector. We therefore assume that hairy BH solutions are stable in the range $0.876\lesssim\,p\,\lesssim 1.143$Kokkotas et al. (2017) and that the Schwarzschild BH solutiona stable for $p\gtrsim 0.876$.
@@ V.1 Horizon radius
Figure 9: BH compactness as function of the BH mass, normalized with $\mu$, in GR and in EW gravity. The vertical dotted line correspond to the critical mass $M_{c}$. The other dashed lines correspond to radially unstable branches. Note that for $M>M_{c}$ EW gravity and GR share the same (radially stable) Schwarzschild solution.
@@ V.2 BH spectroscopy & extra modes
Focusing on the most relevant Schwarzschild case, the perturbation equations (33)–(34) in the presence of matter read
$\delta G_{\mu\nu}+\mu^{2}\delta f_{\mu\nu}=0\,,$ (70)
¯$\bar{\Box}\delta f_{\mu\nu}+2\bar{R}_{\mu\sigma\nu\rho}\delta f^{\sigma\rho}-\mu^{2}\delta f_{\mu\nu}=8\pi T_{\mu\nu}\;.$ (71)
Thus, matter perturbations (e.g., the merger initial conditions), would source $\delta f_{\mu\nu}$, whose spectrum consists of massive spin-2 QNMs, and that in turn sources $\delta g_{\mu\nu}$ through Eq. (70).
@@
Thus, the excitation of massive spin-2 QNMs is suppressed in the GR limit $\mu M\gg 1$, as expected. These modes affect the metric perturbations since, by replacing the above equation into Eq. (70), one obtains
$\delta g_{\mu\nu}\sim\frac{\mu^{2}M^{2}}{1-\mu^{2}M^{2}}T_{\mu\nu}M^{2}\,.$ (73)
This shows that the amplitude of the extra modes in the metric is suppressed also in the opposite limit, $\mu M\ll 1$, while when $\mu M\gg 1$ one recovers the ordinary GR scaling, $\delta g_{\mu\nu}\sim T_{\mu\nu}M^{2}$.
In the interesting intermediate regime, $\mu M\sim 1$, the amplitude of the massive spin-2 QNMs is not suppressed in $\delta g_{\mu\nu}$. In this case, the ringdown signal will contain extra modes that can be studied using the methods recently proposed in Ref. Crescimbeni et al. (2024). We postpone a detailed analysis of this interesting problem to future work.
@@ VI Conclusions
Future investigation will focus on the polar sector, which is expected to display the same qualitative features. Furthermore, it would be interesting to compute the signal emitted by inspiralling point particles, to quantify the amplitude of the extra massive spin-2 modes compared to the ordinary GR ones. This would also allow mapping the theory to the recently proposed theory-agnostic framework for extra non-GR modes in the ringdown Crescimbeni et al. (2024).
@@ Appendix A Background semi-analytical functions
$\tilde{f}(x)=\frac{\tilde{f}_{1}}{1+\frac{\tilde{f}_{2}x}{1+\frac{\tilde{f}_{3}x}{1+\frac{\tilde{f}_{4}x}{1+\ldots}}}}\,,\;\tilde{h}(x)=\frac{b_{1}}{1+\frac{\tilde{h}_{2}x}{1+\frac{\tilde{h}_{3}x}{1+\frac{\tilde{h}_{4}x}{1+\ldots}}}}\,,$ (75)
$\epsilon$ $=(1054-1203p)\left(\frac{3}{1271}+\frac{p}{1529}\right)\,,$
$\tilde{f}_{1}$ $=(1054-1203p)\left(\frac{7}{1746}-\frac{5p}{2421}\right)\,,$
$\tilde{h}_{1}$ $=(1054-1203p)\left(\frac{p}{1465}-\frac{2}{1585}\right)\,,$
$\tilde{f}_{2}$ $=\frac{6p^{2}}{17}+\frac{5p}{6}-\frac{131}{102}\,,$
$\tilde{h}_{2}$ $=\frac{81p^{2}}{242}-\frac{109p}{118}-\frac{16}{89}\,,$
$\tilde{f}_{3}$ $=\frac{\dfrac{9921p^{2}}{31}-385p+\dfrac{4857}{29}}{237-223p}\,,$
$\tilde{h}_{3}$ $=-\frac{2p}{57}+\frac{29}{56}\,,$
$\tilde{f}_{4}$ $=\frac{\frac{9p^{2}}{14}+\dfrac{3149p}{42}-\dfrac{2803}{14}}{237-223p}\,,$
@@ Appendix B Perturbations
are the axial vector tensor harmonics, respectively, and
@@ C.1 Continued fraction method
$\begin{split}&\boldsymbol{\alpha}_{n}=\begin{pmatrix}\alpha_{n}&0\\ 0&\alpha_{n}\end{pmatrix}\,,\;\boldsymbol{\beta}_{n}=\begin{pmatrix}\beta_{n}&\Lambda-2\\ -2&\beta_{n}-3\end{pmatrix}\\[5.69054pt] &\boldsymbol{\gamma}_{n}=\begin{pmatrix}\gamma_{n}&6-3\Lambda\\ 0&\gamma_{n}+9\end{pmatrix}\,,\;\Lambda\equiv\ell(\ell+1)\,,\end{split}$ (88)
@@
$+\frac{\omega^{2}(2n-4i\omega+1)}{q}-3q(2n-4i\omega+1)$ (90)
$+4i(2n+1)\omega-4q^{2}+12\omega^{2}\,,$
$\gamma_{n}=$ $\;q^{-2}[q^{2}\left(n^{2}-4in\omega-6\omega^{2}-9\right)+q^{4}+\omega^{4}$
$+2q^{3}(n-2i\omega)-2q\omega^{2}(n-2i\omega)]\,.$ (91)
@@ C.2 Direct forward integration
each component of which is made of the coefficients $B_{i}^{(j)}$. The superscript $j$ denotes a particular vector of the chosen basis, and corresponds to one of the different perturbation functions. The subscript $i$ characterizes each one of the $N$ independent integrations. The QNM frequency $\omega_{0}=\omega_{R}+i\omega_{I}$ will then correspond to the solutions of

### v3

@@ IV.2 Hairy BH background
From (29) we have three nontrivial components. Below we show their schematic form (note that the coefficients depend on the background coordinates and on $\omega$):
@@
We then solve the constraint equation together with the $(\theta\varphi)$ component of (29) for $h_{0},\,F_{0}$ and we substitute those into the $(r\varphi)$ component of (29) and the $(r\varphi)$ and $(\theta\varphi)$ components of (30). We end up with the following coupled system of equations for the three perturbation functions $h_{1},\,F_{1},\,F_{2}$:
@@
where $\boldsymbol{\Psi}=(\Psi^{(1)},\Psi^{(2)},\Psi^{(3)})\equiv(h_{1},F_{1},F_{2})$,
@@
where $c$ and $b_{1}$ are the coefficients appearing in the near-horizon expansions of the background solutions (15), (16). By substituting in (61) we can solve for $h_{1}^{(n)},\,F_{1,2}^{(n)}$ with $n\geq 1$ in terms of three free parameters, namely $(F_{1}^{(0)},F_{2}^{(0)},h_{1}^{(0)})$.
Figure 5: Real (upper panel) and imaginary (lower panel) parts of the fundamental axial, $\ell=2$ vector modes, as functions of $2M\mu$ in the case of hairy BH background. The solid lines correspond to the frequencies of the massless-led and massive-led modes, computed using the DI approach with $n_{\text{max}}=8$. For comparison, we also show the frequencies of the corresponding modes in the case of Schwarzschild background (dashed lines). The vertical dashed line denotes the threshold $M=M_{c}$.
@@
For the forward integration, the expansions at infinity have to be modified in order to incorporate the contribution from the massive and massless modes as well. For QNMs, we look for solutions with $\omega>\mu$ and impose outgoing boundary conditions at infinity, leading to:
@@
where $k=\sqrt{\omega^{2}-\mu^{2}}$ and $x=M(\mu^{2}-2\omega^{2})/(ik)$. After substitution in (61) and solving order by order we determine all coefficients in terms of three parameters, namely $(H_{1b}^{(0)},F_{1a}^{(0)},F_{2a}^{(0)})$. To find the QNMs we perform three forward integrations by fixing the horizon parameters $(h_{1}^{(0)},F_{1}^{(0)},F_{1}^{(0)})$, and then demand that the coefficients associated with the ingoing waves at infinity vanish, according to the method explained in Appendix C.
Since the massless and massive perturbations are coupled, we do not have purely massive and purely massless QNM solutions. We find, instead, massless-led solutions, in which the contribution of the massive field is due to the coupling in Eq. (6), and similarly massive-led solutions. Therefore, we have four series of modes: the massless-led vector and tensor modes (which in this case are different from those of GR), and the massive-led vector and tensor modes.
@@
For the vector modes showed in Fig. 5 we notice that within the parameter space we explore, both the decay rate and oscillation frequency increase as the deviation from GR increases, i.e. for smaller values of $\mu$ (larger values of the coupling constant $\alpha$) at a fixed BH mass $M$. In the case of the tensor modes depicted in Fig. 6 the behavior of the massive-led mode is non-monotonic and harder to describe. For clarity, in Figs. 7 and 8, we present the same QNM modes as functions of $p=\mu r_{h}$, and normalizing the frequencies and the mass of the spin-2 mode with the BH horizon radius rather than with its mass.
At variance with with the case of the Schwarzschild background discussed in the previous subsection, the hairy background is numerical or semi-analytical, and therefore a CF approach for the calculation of QNMs is not as straightforwardly applicable. Thus, we only employ the DI approach. Due to the limitations of the DI method, we do expect numerical errors to be present in this computation. However, the analysis of the Schwarzschild background case gives strong indication that the DI approach with $n_{\text{max}}=8$ is accurate enough to give the QNM frequencies within a few percent error for the mass range we are interested in.
@@ V Phenomenology
In the region where both Schwarzschild and hairy BHs are radially stable, our analysis did not find any numerical evidence for the instability of either solutions in the axial sector. We therefore assume that hairy BH solutions are stable in the range $0.876\lesssim\,p\,\lesssim 1.143$Kokkotas et al. (2017), and that Schwarzschild BH solutions are stable for $p\gtrsim 0.876$.
@@ V.1 Horizon radius
Figure 9: BH compactness as function of the BH mass, normalized with $\mu$, in GR and in EW gravity. The vertical dotted line correspond to the critical mass $M_{c}$. The other dashed lines correspond, to radially unstable branches. Note that for $M>M_{c}$, EW gravity and GR share the same (radially stable) Schwarzschild solution.
@@ V.2 BH spectroscopy & extra modes
It is also worth noting that the massless modes in the Schwarzschild background coincide with those of GR, and all other modes shown in Section IV do not converge to GR modes as $\alpha\to 0$. Thus, from an EFT perspective there are no corrections to the QNM spectrum at the leading order in $\alpha$, but the new non-GR modes might be excited with amplitudes that vanish in the $\alpha\to 0$ limit.
@@
Focusing on the most relevant Schwarzschild case, the perturbation equations (33)–(34) in the presence of matter read fol
$\delta G_{\mu\nu}+\mu^{2}\delta f_{\mu\nu}=\frac{8\pi}{3}g_{\mu\nu}T\,,$ (70)
¯$\bar{\Box}\delta f_{\mu\nu}+2\bar{R}_{\rho\mu\sigma\nu}\delta f^{\rho\sigma}-\mu^{2}\delta f_{\mu\nu}=$
$\qquad\qquad 8\pi\left(T_{\mu\nu}-\frac{1}{3}g_{\mu\nu}T+\frac{1}{3\mu^{2}}\bar{\nabla}_{\mu}\bar{\nabla}_{\mu}T\right)\;.$ (71)
Thus, matter perturbations (e.g., the merger initial conditions) would source $\delta f_{\mu\nu}$, whose spectrum consists of massive spin-2 QNMs, and that in turn sources $\delta g_{\mu\nu}$ through Eq. (70).
@@
Thus, the excitation of massive spin-2 QNMs is suppressed in the GR limit $\mu M\gg 1$, as expected. These modes affect the metric perturbations since, by replacing the above equation into Eq. (70), in the $\mu M\gg 1$ limit one obtains
$\delta G_{\mu\nu}=8\pi T_{\mu\nu}+{\cal O}(\mu^{-2}M^{-2})$ (73)
and hence one recovers the ordinary GR result.
The above heuristic argument suggests that, in the GR limit, the extra modes will be present in the ordinary gravitational ringdown at ${\cal O}(\mu^{-2})={\cal O}(\alpha)$. However, within the EFT framework, through a field redefinition, it can be shown that that theory in vacuum is equivalent to GR to leading other in the quadratic-curvature couplings Endlich et al. (2017). Thus, an interesting question is whether some cancellation occurs and the new-mode excitation appears at ${\cal O}(\alpha^{2})$, i.e. at the same order of the theories with quartic terms in the curvature tensor, considered e.g. in Endlich et al. (2017); Cardoso et al. (2018). This problem can be addressed by modelling through perturbation theory the actual excitation of the QNMs by a physical source, like e.g. an inspiralling particle.
In the interesting intermediate regime, $\mu M\sim 1$ (which is however not covered by the EFT), the amplitude of the massive spin-2 QNMs is not suppressed in $\delta g_{\mu\nu}$. In this case, the ringdown signal will contain extra modes that can be studied using the methods recently proposed in Ref. Crescimbeni et al. (2024).
We postpone a detailed analysis of these interesting problems, as well as a comparison with the EFT prediction, to an ongoing follow-up work fol .
@@ VI Conclusions
Future investigation will focus on the polar sector, which is expected to display the same qualitative features. Furthermore, it would be interesting to compute the signal emitted by inspiralling point particles, to quantify the amplitude of the extra massive spin-2 modes compared to the ordinary GR ones (see also Ould El Hadj (2024) for some recent relevant results in massive gravity). This would also allow mapping the theory to the recently proposed theory-agnostic framework for extra non-GR modes in the ringdown Crescimbeni et al. (2024). Moreover, this would clarify the actual order of the effects of this theory in case it is considered in an EFT expansion.
@@ Appendix A Background semi-analytical functions
$\tilde{f}(x)=\frac{\tilde{f}_{1}}{1+\frac{\tilde{f}_{2}x}{1+\frac{\tilde{f}_{3}x}{1+\frac{\tilde{f}_{4}x}{1+\ldots}}}}\,,$ (75)
$\;\tilde{h}(x)=\frac{b_{1}}{1+\frac{\tilde{h}_{2}x}{1+\frac{\tilde{h}_{3}x}{1+\frac{\tilde{h}_{4}x}{1+\ldots}}}}\,,$ (76)
$\epsilon$ $=(1054-1203p)\left(\frac{3}{1271}+\frac{p}{1529}\right)\,,$ (77)
$\tilde{f}_{1}$ $=(1054-1203p)\left(\frac{7}{1746}-\frac{5p}{2421}\right)\,,$ (78)
$\tilde{h}_{1}$ $=(1054-1203p)\left(\frac{p}{1465}-\frac{2}{1585}\right)\,,$ (79)
$\tilde{f}_{2}$ $=\frac{6p^{2}}{17}+\frac{5p}{6}-\frac{131}{102}\,,$ (80)
$\tilde{h}_{2}$ $=\frac{81p^{2}}{242}-\frac{109p}{118}-\frac{16}{89}\,,$ (81)
$\tilde{f}_{3}$ $=\frac{\dfrac{9921p^{2}}{31}-385p+\dfrac{4857}{29}}{237-223p}\,,$ (82)
$\tilde{h}_{3}$ $=-\frac{2p}{57}+\frac{29}{56}\,,$ (83)
$\tilde{f}_{4}$ $=\frac{\frac{9p^{2}}{14}+\dfrac{3149p}{42}-\dfrac{2803}{14}}{237-223p}\,,$ (84)
@@ Appendix B Perturbations
are the axial vector harmonics, and
@@ C.1 Continued fraction method
$\boldsymbol{\alpha}_{n}=\begin{pmatrix}\alpha_{n}&0\\ 0&\alpha_{n}\end{pmatrix}\,,$ (97)
$\boldsymbol{\beta}_{n}=\begin{pmatrix}\beta_{n}&\Lambda-2\\ -2&\beta_{n}-3\end{pmatrix}\,,$ (98)
$\boldsymbol{\gamma}_{n}=\begin{pmatrix}\gamma_{n}&6-3\Lambda\\ 0&\gamma_{n}+9\end{pmatrix}\,,$ (99)
$\Lambda\equiv\ell(\ell+1)\,,$ (100)
@@
$+\frac{\omega^{2}(2n-4i\omega+1)}{q}$
$-3q(2n-4i\omega+1)$
$+4i(2n+1)\omega-4q^{2}+12\omega^{2}\,,$ (102)
$\gamma_{n}=$ $\;q^{-2}[q^{2}\left(n^{2}-4in\omega-6\omega^{2}-9\right)+q^{4}$
$+\omega^{4}+2q^{3}(n-2i\omega)-2q\omega^{2}(n-2i\omega)]\,.$ (103)
@@ C.2 Direct forward integration
the components of which are the coefficients $B_{i}^{(j)}$. The superscript $j$ denotes a particular vector of the chosen basis, and corresponds to one of the different perturbation functions. The subscript $i$ characterizes each one of the $N$ independent integrations. The QNM frequency $\omega_{0}=\omega_{R}+i\omega_{I}$ will then correspond to the solutions of